OpenStudy (anonymous):

Teas costing $7 per pound are to be mixed with teas costing $4 per pound to make a 30-pound mixture. If the mixture is to sell for $6 per pound, how many pounds of each kind of tea should be used?

OpenStudy (amistre64):

e+c = 30; e = 30-c e(7) + c(4) = 30(6) (30-c)(7) + c(4) = 180

OpenStudy (amistre64):

a+b = c ; a = c-b ax + by = cz (c-b)x + by = cz cx - bx + by = cz b(y-x) = c(z-x) b = c(z-x)/(y-x) ; i dont think I could remember that to well :)

OpenStudy (amistre64):

cheap stuff = (30)(-1)(-3) = 10 if its right :) 20(7) + 10(4) = 30(6) maybe? 140 + 40 = 180 .. its right